R. Bachimanchi, K. Baggett, A. Freyberger, R. Geng, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, T. Powers, T. Reilly, R. Rimmer, A. Solopova, S. Suhring, and A. Seryi.
Agenda and Notes
1. Operations Status and Summary/Review (Freyberger)
Last two weeks physics running has improved considerably. Regularly have all four halls receiving over 7 hours of beam.
This morning all C100 cavities in the linacs indicating blue on comfort screen.
2. C100 Planning (Solopova)
3. Planning for particulate studies (Geng)
Strong encouragement from the AAC to do serious studies "for the community. Could the LERF be hijacked?
4. Plasma Cleaning (Powers)
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?